OpenStudy (tomektomeknyc):

How to solve it : tan 490 degrees as a function of an angle less than 45 degrees is: a) tan 40 degrees, b) - tan 40 degrees c) cot 40 degrees d) - cot 40 degrees

OpenStudy (ranga):

You can subtract a full 360 degrees out of it without affecting any trigonometric function. tan(490) = tan(490-360) = tan(130) Since all the choices involve 40 degrees, we need to write it as: tan(130) = tan(90 + 40) = - cot(40)